OpenStudy (anonymous):

Solve differential equation y''-k^2*y=0 with initial conditions y'(0)=-y and y'(-d)=0. Thanks

OpenStudy (zzr0ck3r):

can you use leplace yet?

OpenStudy (zzr0ck3r):

or integration method? let u = e^int(-k^2)dt

OpenStudy (unklerhaukus):

\[y''-k^2y=0\qquad\Longrightarrow\qquad y(x)=A\sinh(kx)+B\cosh(kx)\]

OpenStudy (anonymous):

@aliyuaziz you mean y'(0)=-y(0)?

OpenStudy (anonymous):

my issue is with the boundary condition which contains y: y'(0) = -y.

OpenStudy (anonymous):

can't use laplace yet please

OpenStudy (anonymous):

i haven't seen boundary condition like that!!!

OpenStudy (anonymous):

This is my first time to encounter that type of boundary condition too.

OpenStudy (unklerhaukus):

can you find \[y^\prime(x)\]

OpenStudy (anonymous):

y'(x) = Akcosh(kx) + Bksinh(kx)

OpenStudy (unklerhaukus):

now you can substitute those boundary conditions

OpenStudy (anonymous):

i think that is not a correct type of boundary condition!!!

OpenStudy (anonymous):

According to UnkleRhaukus, y(x) = Asinh(kx) + Bcosh(kx) y'(x) = Akcosh(kx) + Bksinh(kx) at y'(0) = -y, A = -y/k and at y'(-d) = 0, B = y/(k*tanh(-kd) therefore y(x) = -y/k*sinh(kx) + y/k*cosh(kx)/tanh(-kd) correct?

OpenStudy (anonymous):

you treated y like a constant that is not correct

OpenStudy (anonymous):

in Ordinary Differential Equations we have not boundary condition like that! could u plz check the question again? maybe it is y'(0)=-y(0)

OpenStudy (anonymous):

or something else

OpenStudy (anonymous):

@mukushla, thanks for pointing out the treatment of y as a constant which is wrong. The issue is on how do we solve solve problems where the boundary condition is not constant but dependent on the solution?

OpenStudy (anonymous):

let me say something

OpenStudy (anonymous):

if y'(0)=-y then y=-(A+B) so y is constant then y'=0 and y''=0 and the equation will be -k^2y=0 so y=0!!!!

OpenStudy (anonymous):

err y=-B